Use the data below for the solubility of 1-butane in benzyl alcohol at 273K to calculate the Henry’s law constant for this gas.

P2/torr X2
200 0.040
400 0.087
600 0.151
700 0.193
760 0.226
Answer #1

Henry law states that the solubility of a gas in a liquid is directly proportional to the partial pressure of a gas over the liquid.

P = KH*X

KH = P/X

KH = henry law constant

P = partial pressure of 1-butane

X = solubility of 1-butane in benzyl alcohol

P /torr X 200 400 600 KH =P/X average 0.04 5000 0.087 4597.701 4112.197 0.151 3973.51 0.193 3626.943 0.226 3362.832 700 760

KH = henry law constant = 4112.2 torr